OpenStudy (anonymous):

Solve by completing the square. x² - 2x - 24 = 0

OpenStudy (anonymous):

x^2-2x=24 x^2-2x+c-c=24 c=[(-2)/2]^2=(-1)^2=1 x^2-2x+1-1=24 (x-1)^2-1=24 (x-1)^2=25 sqrt[(x-1)^2]=sqrt(25) x-1=positive or negative 5= +/-5 x-1=+/-5 x=1+/-5 x=6, x=-4